OpenStudy (anonymous):

I need to find the surface area and volume of a sphere. I must solve it completely. r=x in the problem. How am I suposed to solve that?

OpenStudy (anonymous):

surface area=\[4\times \pi \times x^{2}\] volume=\[4/3\times \pi \times x^3\]